# CohEx: A Generalized Framework for Cohort Explanation

### v1

@@ Abstract
eXplainable Artificial Intelligence (XAI) has garnered significant attention for enhancing transparency and trust in machine learning models. However, the scopes of most existing explanation techniques focus either on offering a holistic view of the explainee model (global explanation) or on individual instances (local explanation), while the middle ground, i.e., cohort-based explanation, is less explored. Cohort explanations offer insights into the explainee’s behavior on a specific group or cohort of instances, enabling a deeper understanding of model decisions within a defined context. In this paper, we discuss the unique challenges and opportunities associated with measuring cohort explanations, define their desired properties, and create a generalized framework for generating cohort explanations based on supervised clustering.
@@ 1 Introduction & Motivation
eXplainable Artificial Intelligence aims to improve the existing machine learning system by providing transparent explanations for model predictions and decisions (Arrieta et al. 2020; Adadi and Berrada 2018). The scope of explanation refers to the degree to which the explanation provided is generalized. The scope, as defined by (Sokol and Flach 2020), is usually (1) global explanation, which aims to explain the explainee comprehensively; or (2) local explanation, which focuses on explaining the behavior in one instance of model prediction. However, there is a less explored middle ground, namely (3) cohort explanation, or sometimes also called regional explanation, which generalizes the explanation on a subset of a dataset, a subspace in the model’s input space or decision space. Cohort explanation can be seen as a spectrum that generalizes both local and global explanation methods, where global and local explanations are at the two extremes of conciseness and faithfulness. This generalization allows users to uncover patterns, biases, and context-specific insights that may not be apparent when examining individual instances in isolation or the model as a whole.
(a) Data and the explainee decision tree.
@@
Figure 1: The motivating example: a synthesized patient binary classification problem. (b, top) Global importance; (b, bottom) Local importance; (c, top) Cohort importance on young patients with high family history; (c, middle) Cohort importance on middle-aged patients or young patients with low family history; (d, bottom) Cohort explanation on older patients.
@@
To illustrate this motivation, consider a simplified medical patient classification task in Fig. 1(a). The two features are age and the fraction of family members with this disease. The likelihood of the condition occurring increases with respect to both features. The explainee model is a decision tree with a depth of two. Note that we treat the decision tree as a black-box, i.e. the explainer does not know that the model is a tree.
@@ 3 Notation
  • $M$: explainee model, or the model to be explained;
@@
  • $\omega$: a local explanation method. It is data driven and outputs local feature importance. Specifically, $\omega_{M}(X,x)$ is the result of evaluating $\omega$ in $x$, on the model $M$ and using $X$ as context samples. It should treat the explainee $M$ as a black-box, i.e., $\omega$ does not depend on the specific architecture of $M$;
@@ 4 Desiderata
Stability w.r.t. explainee model
w.r.t explainee model
@@
where $e_{j}$ is evaluated on explainee $M$ and $\tilde{e}_{j}$ is on $\tilde{M}_{j}$, and the expectation is over the randomness from $p$. An ideal cohort explanation should have a low locality loss and thus focus on localized information.
@@
The explanation should be robust w.r.t. (1) the explainee model and (2) the cohort definitions. The cohort definitions should be stable and the importance should not drastically differ when the definitions are slightly changed.
@@
To evaluate importance stability, we consider the change in explanation in a particular cohort if a new random sample is added to this cohort. Namely, we evaluate importance stability as
@@ 5 Problem Definition
We emphasize on generate post-hoc cohort explanation through existing local, data driven, feature importance explanations. We formulate the cohort explanation as an optimization problem, with the goal of minimizing the generalizability loss in Eq. 1 while penalizing for having too many cohorts. Consider a local data-driven explanation method $\omega_{M}(X,x)$, where $M$ is the explainee model, $X$ is a context dataset and $x$ is a sample. The problem is defined as
@@
Input : Number of iterations $n$, Expected number of cohorts $k$, Explainee model $M$,
@@ 6 Challenges
Figure 2: Two regression models with the same behavior on a specific cohort. The orange circle represents the cohort.
@@ 7 Approach
To generate the cohort definitions, we choose supervised clustering. Compared to traditional clustering, it assumes that data are labeled and uses both features $X$ and labels $y$ to group data. Each data sample is clustered into the closest centroids in the feature space, while the goodness of an assignment is based on the labels, or in our case, feature importance (Eick et al., 2004). This ensures that the cohort definitions are always closed and well defined, while constraining local explanations within the cohort to be similar for the generalizability criteria. Specifically, in our evaluations, we use Single Representative Insertion/Deletion Steepest Decent Hill Climbing with Randomized Restart (SRIDHCR) as the supervised learning method, a greedy algorithm that attempts to solve the problem by randomly adding or removing centroids (Eick, Zeidat, and Zhao 2004).
@@ 8 Evaluation
  • REPID, a tree-based partitioning method (Herbinger, Bischl, and Casalicchio 2022). For the sake of comparison, we modify the algorithm to use a local explainer that is consistent with other method, instead of the original partial dependency explainer;
@@ 8.2 Bike Sharing
This experiment focuses on a realistic machine learning regression problem. The bike sharing task asks to predict the hourly number of rented bikes between years 2011 and 2012 in the Capital bikeshare system based on seasonal and weather information. For the explainee model, we train an XGBoost model (Chen and Guestrin 2016) on this dataset and achieve a $R^{2}$ score of 0.9453. We continue to use $k^{*}=4$ or a depth of two for a similar number of cohorts and use SHAP as the local explainer.
@@ 8.3 Digit Classification
We evaluate based on MNIST (Deng 2012). Note that the MNIST inputs are cropped, centered, and relatively simple, so measuring the distance between importance heatmaps is meaningful. We built the explainee model as a neural network with two convolutional layers and two fully-connected layers and achieved an accuracy of 93.93% on the test dataset. More details on the model and a visualization of the cohort explanation can be found in Appendix. C.3.
@@ 8.4 Evaluation Metrics and Analysis
Generalizability & Conciseness CohEx achieves the best performance based on the objective function in Eq. 6. Note that all methods yield $k=4$, so the conciseness loss for all is zero. VINE yields a low generalizability loss since its clustering criterion focuses only on feature importance, while all other methods need to consider the distance in the feature space. However, CohEx is able to exceed VINE since they use different local importance scores $w_{x}$. Through the recomputing of importance, CohEx edits the importance by putting more emphasis on local structures, allowing the importance scores in each cohort to have lower variance. Introducing GALE re-weighting improves the performance of VINE and REPID since it also recomputes importance, but it does not perform well on low-dimensional features as shown in Fig. 3 since the importance of the feature with the lowest homogeneity will always be set to zero.
@@
Cohort Locality CohEx yields a significantly lower locality loss compared to the other methods according to Eq. 3, which is as expected since CohEx introduces an iterative refinement process specifically for locality. To further justify the approach, in the patient classification problem, CohEx identifies a clustering that mostly match with the three clusters in our ideal explanation proposed in Sect. 1. In particular, CohEx is the only algorithm that identifies a region (the blue cohort in Fig. 3(f)) in which family history is the more important feature. Note that for almost all samples, the static LIME importance of age is higher than family history similar to the result in Fig. 1(b). Therefore, since other methods are based on a static set of local importance scores, they cannot find any regions with a meaningful family history importance. We further justify that such a region should exist, since if we only consider said region, the horizontal decision boundary is the dominant decision boundary. This shows that CohEx’s explanation is more local to its cohorts. We also note that applying GALE would harm locality, since the re-weighting process smoothes the importance over the whole dataset, specifically contradicts the goal of locality.
@@
Disjoint Cohorts Only VINE fails this property since it clusters solely based on importance and it’s very likely that some distant samples happen to have similar explanations and thus are partitioned together. All other method guarantees that the cohorts can be separated into disjoint areas by hyperplanes.
@@
Cohort Stability In all scenarios, REPID achieves higher ARI scores based on Eq. 4, which shows that our supervised clustering-based approaches introduce variance. We consider this variance to be inherent to the methods, as both the recomputing process and SRIDHCR are trial-and-error by nature. The question of how to improve the stability remains a challenge for future work.
@@
Importance Stability It measures the change in cohort explanations, if a random sample is added to the cohort. For fair comparisons, we allow the baseline methods to recompute the importances similar to CohEx after adding the sample, which greatly improve their stabilities. CohEx still exceeds the baselines. This demonstrates that even though the partitions of CohEx is not as robust based on the previous cohort stability metric, on a fixed cohort, CohEx’s explanation is much less gullible to changes or noises in the dataset.
@@ Appendix B CohEx w/ Non-Data-Driven Methods
Although we focus on adapting data-driven methods, CohEx can be modified if any of these premises cannot be met. If the base method is not data driven, for example, saliency map (saliency), vanilla gradient, SmoothGrad (smoothgrad) or Grad×$\times$×Input (gradxinput), the explanations will not change during recomputing since $w$ does not depend on the input dataset $X$. In this case, similar to cohort explanations, we only need to run one pass of lines 6-13 in Algorithm 1. The algorithm fallbacks to a hierarchical scheme that first computes the importances, then uses supervised clustering to create the cohort definitions. However, as discussed previously, without the iterative loop, the result may not satisfy the cohort locality property.
@@ C.3 Digit Classification
We built the explainee model as a simple convolutional neural network with two convolutional layers followed by two fully connect layers. The convolutional layers each have a kernel size of 5, and maps to 10 and 20 channels, respectively. Both layers are followed by a max-pooling layer and then a ReLU activation, while the second layer has an additional dropout layer between the convolutional and the pooling layers. The fully connected layers map the final $20\times 4\times 4$ pixels to 50 features, then a ReLU activation and a dropout are maps to ten class outputs by another fully connected layer. We trained the model for two epochs with an SGD optimizer and achieved an accuracy of 93.93% on the test dataset. The model is trained using the full training dataset with all digits.
To demonstrate the properties of cohort explanations, we only include a small set of 200 images of digits 7 and 9 during explanation evaluation. This experiment aims to see if CohEx can focus on pixels that separate these specific digits. Note that the network is trained using the full 10-digits and outputs a probability for each of the classes. We opt to use DeepLIFT (deeplift) as the base explainer, an approximation of SHAP on deep models. DeepLIFT will generate ten feature importance images with respect to all ten classes, even if only two classes are present in the testing dataset. We set the expected number of cohorts to 4, allowing the framework to detect interesting subgroups within each digit class.

### v2

@@ Abstract
eXplainable Artificial Intelligence (XAI) has garnered significant attention for enhancing transparency and trust in machine learning models. However, the scopes of most existing explanation techniques focus either on offering a holistic view of the target model (global explanation) or on individual instances (local explanation), while the middle ground, i.e., cohort-based explanation, is less explored. Cohort explanations offer insights into the target model’s behavior on a specific group or cohort of instances, enabling a deeper understanding of model decisions within a defined context. In this paper, we discuss the unique challenges and opportunities associated with measuring cohort explanations, define their desired properties, and create a generalized framework for generating cohort explanations based on supervised clustering.
Codehttps://github.com/fy-meng/cohex
@@ 1 Introduction & Motivation
eXplainable Artificial Intelligence aims to improve the existing machine learning system by providing transparent explanations for model predictions and decisions (Arrieta et al. 2020; Adadi and Berrada 2018). The scope of explanation refers to the degree to which the explanation provided is generalized. The scope, as defined by (Sokol and Flach 2020), is usually (1) global explanation, which aims to explain the target model comprehensively; or (2) local explanation, which focuses on explaining the behavior in one instance of model prediction. However, there is a less explored middle ground, namely (3) cohort explanation, or sometimes also called regional explanation, which generalizes the explanation on a subset of a dataset, a subspace in the model’s input space or decision space. Cohort explanation can be seen as a spectrum that generalizes both local and global explanation methods, where global and local explanations are at the two extremes of conciseness and faithfulness. This generalization allows users to uncover patterns, biases, and context-specific insights that may not be apparent when examining individual instances in isolation or the model as a whole.
(a) Data and the target decision tree.
@@
Figure 1: A motivating example: a synthesized patient binary classification problem.
Lack of Robustness of Feature Attribution Methods
Many existing explanation methods in XAI, notably feature attribution methods, face limitations regarding robustness. The output can be sensitive to small perturbations in input or model parameters, leading to unstable or unreliable insights (Alvarez-Melis and Jaakkola 2018; Slack et al. 2020). Some explainer would be produce biased output under simple transformations, failing to only capture relevant information for decision-making (Kindermans et al. 2019). A common practice to improve robustness in machine learning is via bootstrapping - averaging the output of multiple weak learners to reduce overfitting (Breiman 1996, 2001; Sagi and Rokach 2018). A similar approach could also be beneficial for XAI. By averaging the explanations on a group of similar samples, we could reduce the noises in explanations, providing more stable, interpretable insights that account for patterns across relevant subsets of data.
@@
To illustrate this motivation, consider a simplified medical patient classification task in Fig. 1(a). The two features are age and the fraction of family members with this disease. The likelihood of the condition occurring increases with respect to both features. The target model is a decision tree with a depth of two. Note that we treat the decision tree as a black-box, i.e. the explainer does not know that the model is a tree.
@@ 3 Notation
  • $M$: target model, or the model to be explained;
@@
  • $\omega$: a local explanation method. It is data driven and outputs local feature importance. Specifically, $\omega_{M}(X,x)$ is the result of evaluating $\omega$ in $x$, on the model $M$ and using $X$ as context samples. It should treat the target $M$ as a black-box, i.e., $\omega$ does not depend on the specific architecture of $M$;
@@ 4 Desiderata
Stability w.r.t. target model
w.r.t target model
@@
where $e_{j}$ is evaluated on target $M$ and $\tilde{e}_{j}$ is on $\tilde{M}_{j}$, and the expectation is over the randomness from $p$. An ideal cohort explanation should have a low locality loss and thus focus on localized information.
@@
The explanation should be robust w.r.t. (1) the target model and (2) the cohort definitions. The cohort definitions should be stable and the importance should not drastically differ when the definitions are slightly changed.
@@
To evaluate importance stability, we consider the change in explanation in a particular cohort if a new random sample is added to this cohort. Ideally, we want all samples to have localized, stable local importance scores. Thus, adding a sample to a cohort should not drastically affect its average importance. Therefore, we evaluate importance stability as
@@ 5 Problem Definition
We emphasize on generate post-hoc cohort explanation through existing local, data driven, feature importance explanations. We formulate the cohort explanation as an optimization problem, with the goal of minimizing the generalizability loss in Eq. 1 while penalizing for having too many cohorts. Consider a local data-driven explanation method $\omega_{M}(X,x)$, where $M$ is the target model, $X$ is a context dataset and $x$ is a sample. The problem is defined as
@@
Input : Number of iterations $n$, Expected number of cohorts $k$, Target model $M$,
@@ 6 Challenges
Figure 2: Two 1-D regression models with the same behavior on a specific cohort. The horizontal axis represents the feature, and the vertical axis represents the model output. The orange circle denotes the cohort.
@@ 7 Approach
To generate the cohort definitions, we choose supervised clustering. Compared to traditional clustering, it assumes that data are labeled and uses both features $X$ and labels $y$ to group data. Each data sample is clustered into the closest centroids in the feature space, while the goodness of an assignment is based on the labels, or in our case, feature importance (Eick et al., 2004). This ensures that the cohort definitions are always closed and well defined, while constraining local explanations within the cohort to be similar for the generalizability criteria. Specifically, in our evaluations, we use Single Representative Insertion/Deletion Steepest Decent Hill Climbing with Randomized Restart (SRIDHCR) as the supervised learning method, a greedy algorithm that attempts to solve the problem by randomly adding or removing centroids (Eick, Zeidat, and Zhao 2004). Note that standard, unsupervised clustering algorithms such as HDBSCAN does not apply here since we need a supervised mechanism to cluster both the feature values and the importance scores simultaneously.
@@ 8 Evaluation
  • REPID, a tree-based partitioning method (Herbinger, Bischl, and Casalicchio 2022). For the sake of comparison, we modify the algorithm to use a local explainer that is consistent with other method, instead of the original partial dependency explainer;  [^1] Python’s SHAP package implements an auto-cohort feature. Though no reference is provided, we believe that it is equivalent to REPID.
@@ 8.2 Bike Sharing
This experiment focuses on a realistic machine learning regression problem. The bike sharing task asks to predict the hourly number of rented bikes between years 2011 and 2012 in the Capital bikeshare system based on seasonal and weather information. For the target model, we train an XGBoost model (Chen and Guestrin 2016) on this dataset and achieve a $R^{2}$ score of 0.9453. We continue to use $k^{*}=4$ or a depth of two for a similar number of cohorts and use SHAP as the local explainer.
@@ 8.3 Digit Classification
We evaluate based on MNIST (Deng 2012). Note that the MNIST inputs are cropped, centered, and relatively simple, so measuring the distance between importance heatmaps is meaningful. We built the target model as a neural network with two convolutional layers and two fully-connected layers and achieved an accuracy of 93.93% on the test dataset. More details on the model and a visualization of the cohort explanation can be found in Appendix. C.3.
@@ 8.4 Evaluation Metrics and Analysis
CohEx achieves the best performance based on the objective function in Eq. 6. Note that all methods yield $k=4$, so the conciseness loss for all is zero. VINE yields a low generalizability loss since its clustering criterion focuses only on feature importance, while all other methods need to consider the distance in the feature space. However, CohEx is able to exceed VINE since they use different local importance scores $w_{x}$. Through the recomputing of importance, CohEx edits the importance by putting more emphasis on local structures, allowing the importance scores in each cohort to have lower variance. Introducing GALE re-weighting improves the performance of VINE and REPID since it also recomputes importance, but it does not perform well on low-dimensional features as shown in Fig. 3 since the importance of the feature with the lowest homogeneity will always be set to zero.
Cohort Locality
@@
CohEx yields a significantly lower locality loss compared to the other methods according to Eq. 3, which is as expected since CohEx introduces an iterative refinement process specifically for locality. To further justify the approach, in the patient classification problem, CohEx identifies a clustering that mostly match with the three clusters in our ideal explanation proposed in Sect. 1. In particular, CohEx is the only algorithm that identifies a region (the blue cohort in Fig. 3(f)) in which family history is the more important feature. Note that for almost all samples, the static LIME importance of age is higher than family history similar to the result in Fig. 1(b). Therefore, since other methods are based on a static set of local importance scores, they cannot find any regions with a meaningful family history importance. We further justify that such a region should exist, since if we only consider said region, the horizontal decision boundary is the dominant decision boundary. This shows that CohEx’s explanation is more local to its cohorts. We also note that applying GALE would harm locality, since the re-weighting process smoothes the importance over the whole dataset, specifically contradicts the goal of locality.
Disjoint Cohorts
@@
Only VINE fails this property since it clusters solely based on importance and it’s very likely that some distant samples happen to have similar explanations and thus are partitioned together. All other method guarantees that the cohorts can be separated into disjoint areas by hyperplanes.
Cohort Stability
@@
In all scenarios, REPID achieves higher ARI scores based on Eq. 4, which shows that our supervised clustering-based approaches introduce variance. We consider this variance to be inherent to the methods, as both the recomputing process and SRIDHCR are trial-and-error by nature. The question of how to improve the stability remains a challenge for future work.
Importance Stability
@@
It measures the change in cohort explanations, if a random sample is added to the cohort. For fair comparisons, we allow the baseline methods to recompute the importances similar to CohEx after adding the sample, which greatly improve their stabilities. CohEx still exceeds the baselines. This demonstrates that even though the partitions of CohEx is not as robust based on the previous cohort stability metric, on a fixed cohort, CohEx’s explanation is much less gullible to changes or noises in the dataset.
@@ Appendix B CohEx w/ Non-Data-Driven Methods
Although we focus on adapting data-driven methods, CohEx can be modified if any of these premises cannot be met. If the base method is not data driven, for example, saliency map (Simonyan, Vedaldi, and Zisserman 2013), vanilla gradient, SmoothGrad (Smilkov et al. 2017) or Grad×$\times$×Input (Baehrens et al. 2010), the explanations will not change during recomputing since $w$ does not depend on the input dataset $X$. In this case, similar to cohort explanations, we only need to run one pass of lines 6-13 in Algorithm 1. The algorithm fallbacks to a hierarchical scheme that first computes the importances, then uses supervised clustering to create the cohort definitions. However, as discussed previously, without the iterative loop, the result may not satisfy the cohort locality property.
@@ C.3 Digit Classification
We built the target model as a simple convolutional neural network with two convolutional layers followed by two fully connect layers. The convolutional layers each have a kernel size of 5, and maps to 10 and 20 channels, respectively. Both layers are followed by a max-pooling layer and then a ReLU activation, while the second layer has an additional dropout layer between the convolutional and the pooling layers. The fully connected layers map the final $20\times 4\times 4$ pixels to 50 features, then a ReLU activation and a dropout are maps to ten class outputs by another fully connected layer. We trained the model for two epochs with an SGD optimizer and achieved an accuracy of 93.93% on the test dataset. The model is trained using the full training dataset with all digits.
To demonstrate the properties of cohort explanations, we only include a small set of 200 images of digits 7 and 9 during explanation evaluation. This experiment aims to see if CohEx can focus on pixels that separate these specific digits. Note that the network is trained using the full 10-digits and outputs a probability for each of the classes. We opt to use DeepLIFT (Li et al. 2021) as the base explainer, an approximation of SHAP on deep models. DeepLIFT will generate ten feature importance images with respect to all ten classes, even if only two classes are present in the testing dataset. We set the expected number of cohorts to 4, allowing the framework to detect interesting subgroups within each digit class.